# Optimizing Learning for Robust Hyperbolic Deep Learning in Computer Vision

### v2

@@ 4.2 Standard Classification Problem
We follow the experimental setup and hyperparameters of Bdeir et al. (2024) and rely on three main datasets: CIFAR10, CIFAR100, and Mini-Imagenet. We denote HCNN+ as the fully hyperbolic model from Bdeir et al. (2024), updated with efficient convolution implementations and our proposed scaling function. This experiment aims to boost model efficiency and address the performance inconsistencies observed in the original version.
For ResNet-18, we employ HECNN+, following the structure outlined in Bdeir et al. (2024) with alternating Euclidean and hyperbolic blocks, and integrate our efficient convolution operation. While we do not expect similar efficiency gains to those seen in fully hyperbolic models, our focus here is on enhancing scalability. In the case of ResNet-50, we adapt HECNN+ by replacing all blocks with the Lorentz-Core bottleneck block, where we anticipate the most significant efficiency improvements due to the larger model size. Our goal is to evaluate if the Lorentz-Core block can maintain strong performance under these conditions.
For both ResNet-18 and ResNet-50, we use Riemannian SGD with our improved learning scheme and curvature learning. We also decouple the encoder and decoder manifolds, allowing each to independently learn its own curvature for enhanced model flexibility.
@@ Results
For the ResNet-18 experiments, Table 2 shows that the new architectures achieve comparable or better results across almost all cases. The smallest improvements are observed with the hybrid models, likely due to the reduced impact of the hyperbolic components compared to a fully hyperbolic model. However, we see a significant performance gap between the fully hyperbolic models, where our proposed architecture now matches the performance of the hybrid encoders. We hypothesize that the improved scaling helps mitigate the previous performance inconsistencies, with the whole model being two times more memory efficient.
Table 2: Classification accuracy (%) of ResNet-18 models. We estimate the mean and standard deviation from five runs. The best performance is highlighted in bold (higher is better).
CIFAR-10 CIFAR-100 VRAM/tepoch Tiny-ImageNet $(\delta_{rel}=0.26)$ $(\delta_{rel}=0.23)$ For Cifar100 $(\delta_{rel}=0.20)$ Euclidean (He et al., 2016) ${95.14_{\pm 0.12}}$ $77.72_{\pm 0.15}$ 1.2GB - 12s $65.19_{\pm 0.12}$ Hybrid Poincaré (Guo et al., 2022) $95.04_{\pm 0.13}$ $77.19_{\pm 0.50}$ $64.93_{\pm 0.38}$ Hybrid Lorentz (Bdeir et al., 2024) ${94.98_{\pm 0.12}}$ ${78.03_{\pm 0.21}}$ ${65.63_{\pm 0.10}}$ Poincaré ResNet (van Spengler et al., 2023) $94.51_{\pm 0.15}$ $76.60_{\pm 0.32}$ $62.01_{\pm 0.56}$ HECNN Lorentz (Bdeir et al., 2024) ${95.16_{\pm 0.11}}$ ${78.76_{\pm 0.24}}$ 4.3GB - 100s ${65.96_{\pm 0.18}}$ HECNN+ (ours) ${95.15_{\pm 0.07}}$ ${78.80_{\pm 0.12}}$ 3GB - 80s ${65.98_{\pm 0.11}}$ HCNN Lorentz (Bdeir et al., 2024) ${95.15_{\pm 0.08}}$ ${78.07_{\pm 0.17}}$ 10GB - 175s ${65.71_{\pm 0.13}}$ HCNN+ (ours) ${95.17_{\pm 0.09}}$ $\bm{78.81_{\pm 0.19}}$ 5GB - 140s $\bm{66.12_{\pm 0.14}}$
In the ResNet-50 experiments in Table 3, we observe that HECNN+ significantly outperforms the Euclidean model across all datasets, including Tiny-ImageNet, where other models typically begin to show a drop in accuracy. This improvement is likely due to the tighter integration of hyperbolic components and the enhanced scaling, which helps manage the challenges of higher-dimensional embeddings.
We evaluate the impact of our efficient convolution and Lorentz-Core block in Table 3. We see a $\sim 48\%$ reduction in memory usage and $\sim 66\%$ reduction in runtime. We attribute this improvement to the efficient closed-source convolution operations we can now leverage. Similar gains are also observed in generation tasks and classification with smaller models. However, there remains significant room for further optimization when compared to the Euclidean baseline. A key remaining performance bottleneck identified is the batch normalization step, which contributes approximately $60\%$ of the runtime and around $30\%$ of the memory usage. The next step would be to factorize the extensive parallel transports and tangent mappings involved in this operation, potentially alleviating the associated overhead.
Table 3: Classification accuracy (%) of ResNet-50 models. The best performance is highlighted in bold (higher is better).
CIFAR-10 CIFAR-100 VRAM/tepoch Tiny-ImageNet $(\delta_{rel}=0.26)$ $(\delta_{rel}=0.23)$ $(\delta_{rel}=0.20)$ Euclidean (He et al., 2016) 95.1495.14{95.14}95.14 $78.52$ 4.5GB - 30s $66.23$ Hybrid Lorentz (Bdeir et al., 2024) 95.3895.38{95.38}95.38 ${79.35}$ 66.0166.01{66.01}66.01 HECNN (Bdeir et al., 2024) 95.4295.42{95.42}95.42 79.8379.83{79.83}79.83 15.6GB - 300s $66.30$ HCNN+ w Bottleneck Conv. (ours) $\bm{95.46}$ $\bm{80.86}$ 8.1GB - 100s $\bm{67.18}$
@@ 4.3 VAE Image Generation
Previous research has highlighted the ability of hyperbolic neural networks to perform better when using lower dimensional embeddings due to the more expressive space (Nagano et al., 2019; Mathieu et al., 2019; Ovinnikov, 2019; Hsu et al., 2020). A natural extension to this would be implementing the models for VAEs which rely on smaller latent embedding dimensions to encode the inputs. Bdeir et al. (2024) perform this study on image generation and reconstruction for their fully hyperbolic models. In the following section, we reproduce the experimental setup and re-implement the fully hyperbolic VAE using our new efficient convolution and transpose convolution layers. We also use curvature learning with our adjusted Riemannian SGD learning scheme.
@@ Results
We see in table 4 that our implementation of the fully hyperbolic VAE achieves better performance on both datasets. It should also be noted that this is achieved with around 2.5x less memory and at 3x greater training speed. This helps demonstrate both the effectiveness of our proposed curvature learning process and our efficient model components.
@@
CIFAR-100 CelebA Rec. FID Gen. FID Rec. FID Gen. FID Euclidean $63.81_{\pm 0.47}$ $103.54_{\pm 0.84}$ $54.80_{\pm 0.29}$ $79.25_{\pm 0.89}$ Hybrid Poincaré (Mathieu et al., 2019) $62.64_{\pm 0.43}$ $98.19_{\pm 0.57}$ $54.62_{\pm 0.61}$ $81.30_{\pm 0.56}$ Hybrid Lorentz (Nagano et al., 2019) $62.14_{\pm 0.35}$ $98.34_{\pm 0.62}$ $54.64_{\pm 0.34}$ $82.78_{\pm 0.93}$ HCNN Lorentz (Bdeir et al., 2024) $61.44_{\pm 0.64}$ $100.27_{\pm 0.84}$ $54.17_{\pm 0.66}$ $78.11_{\pm 0.95}$ HCNN+ Lorentz (Ours) $57.69_{\pm 0.52}$ $98.14_{\pm 0.44}$ $52.73_{\pm 0.27}$ $77.98_{\pm 0.32}$
@@ 5 Conclusion
In our work, we present many new components and schemas for the use of hyperbolic deep learning in hyperbolic vision. We test these components in three vision tasks and prove the potential of these new components even in float16 conditions.
However, there is still significant room for improvement. Further optimizations to the batch normalization layers could enhance the efficiency of hyperbolic models. Additionally, a key challenge remains with the hyperbolic linear layers when reducing dimensionality. Currently, we match norms to simulate a rotation operation, but we encourage exploring alternative approaches that align more naturally with the mathematical properties of the manifold.
Appendix A Appendix
A.1 Operations in hyperbolic geometry
@@
Each Lorentz transformation can be decomposed via polar decomposition into a Lorentz rotation and a Lorentz boost, expressed as $\mathrm{\textbf{A}}=\mathrm{\textbf{R}}\mathrm{\textbf{B}}$ Moretti (2002). The rotation matrix R is designed to rotate points around the time axis and is defined as
@@
A.2 Scaling Lorentzian Vectors
@@
A.3 Ablations
We test the effect of individual model components in table 5. Each subsequent model involves the default architecture presented in the experimental setup minus the mentioned component. As we can see, the best results are achieved when all the architectural components are included. In the case of attempting to learn the curvature without the proposed optimizer schema, the model breaks completely down due to excessive numerical inaccuracies. One other benefit that we find from learning the curvature is quicker convergence. The model is able to reach convergence in 130 epochs vs the 200 epochs required by a static curve model.

### v3

@@ 4.3 Standard Image Classification Problem
We follow the experimental setup and hyperparameters of Bdeir et al. (2024) and rely on three main datasets: CIFAR10, CIFAR100, and Mini-Imagenet. For ResNet-18, we employ HECNN+, following the structure outlined in Bdeir et al. (2024) with alternating Euclidean and hyperbolic blocks, and integrate our efficient convolution operation. We decouple the encoder and decoder manifolds, allowing each to independently learn its own curvature for enhanced model flexibility. We also implement ResNet-50 models of the proposed hyperbolic hybrid architectures using similarly constructed bottleneck blocks.
@@ Results
For the ResNet-18 experiments, Table 3 shows that the new models are able to remain stable while learning the curvature. Additionally, we see significant performance improvements between HCNN and HCNN+, where our proposed architecture now matches the performance of the hybrid encoders. We hypothesize that the improved scaling function helps mitigate the previous performance inconsistencies. However, the performance difference between the hybrid models is not significant, we hypothesize this is due to the alternating architecture which could limit the effect and instability of hyperbolic components. Finally, both models manage to maintain performance while reducing the memory footprint and runtime by approximately $25-50\%$ and $18-25\%$ respectively. One other benefit that we find from learning the curvature is quicker convergence where the model is able to reach convergence in 130 epochs vs the 200 epochs required by a static curve model.
Table 3: Runtime Analysis for the ResNet-18 models. We report classification accuracy (%) and estimate the mean and standard deviation from five runs. The best performance is highlighted in bold (higher is better).
CIFAR-100 VRAM/tepoch Tiny-ImageNet $(\delta_{rel}=0.23)$ For Cifar100 $(\delta_{rel}=0.20)$ Euclidean $77.72_{\pm 0.15}$ 1.2GB - 12s $65.19_{\pm 0.12}$ Hybrid ($\mathbb{P}$) $77.19_{\pm 0.50}$ $64.93_{\pm 0.38}$ Hybrid ($\mathbb{L}$) ${78.03_{\pm 0.21}}$ ${65.63_{\pm 0.10}}$ ResNet ($\mathbb{P}$) $76.60_{\pm 0.32}$ $62.01_{\pm 0.56}$ HECNN ($\mathbb{L}$) ${78.76_{\pm 0.24}}$ 4.3GB - 100s ${65.96_{\pm 0.18}}$ HECNN+ ${78.80_{\pm 0.12}}$ 3GB - 80s ${65.98_{\pm 0.11}}$ HCNN ($\mathbb{L}$) ${78.07_{\pm 0.17}}$ 10GB - 175s ${65.71_{\pm 0.13}}$ HCNN+ $\bm{78.81_{\pm 0.19}}$ 5GB - 140s $\bm{66.12_{\pm 0.14}}$
In the ResNet-50 experiments in Table 4, HECNN+ significantly outperforms both the Euclidean model and the base hybrid model across both datasets. We also see a $\sim 48\%$ reduction in memory usage and $\sim 66\%$ reduction in runtime. We attribute this improvement to efficient closed-source convolution operations we can now leverage.
Table 4: Runtime Analysis for ResNet-50 models. We report classification accuracy (%) and the best performance is highlighted in bold (higher is better).
CIFAR-100 VRAM/tepoch Tiny-ImageNet $(\delta_{rel}=0.23)$ $(\delta_{rel}=0.20)$ Euclid $78.52$ 4.5GB - 30s $66.23$ HECNN 79.8379.83{79.83}79.83 15.6GB - 300s $66.30$ HECNN+ $\bm{80.86}$ 8.1GB - 100s $\bm{67.18}$
@@ 4.4 VAE Image Generation
Problem Setting and Reference Model
Previous research has highlighted the ability of hyperbolic neural networks to perform better when using lower dimensional embeddings due to the more expressive space (Nagano et al., 2019; Mathieu et al., 2019; Ovinnikov, 2019; Hsu et al., 2020). A natural extension to this would be implementing the models for VAEs which rely on smaller latent embedding dimensions to encode the inputs. Bdeir et al. (2024) perform this study on image generation and reconstruction for their fully hyperbolic models.
@@
We reproduce the experimental setup from (Bdeir et al., 2024) and re-implement the fully hyperbolic VAE using our new efficient convolution and transpose convolution layers. We also use curvature learning with our adjusted Riemannian SGD learning scheme.
@@ Results
Our fully hyperbolic VAE implementation outperforms on both datasets, as shown in Table 5, while using 2.5x less memory and training 3x faster. This highlights the effectiveness of our curvature learning process and efficient model components.
@@
CIFAR-100 CelebA
Rec. FID Gen. FID Rec. FID Gen. FID
Euclid $63.81_{\pm 0.47}$ $103.54_{\pm 0.84}$ $54.80_{\pm 0.29}$ $79.25_{\pm 0.89}$
Hybrid ($\mathbb{P}$) $62.64_{\pm 0.43}$ $98.19_{\pm 0.57}$ $54.62_{\pm 0.61}$ $81.30_{\pm 0.56}$
Hybrid ($\mathbb{L}$) $62.14_{\pm 0.35}$ $98.34_{\pm 0.62}$ $54.64_{\pm 0.34}$ $82.78_{\pm 0.93}$
HCNN ($\mathbb{L}$) $\underline{61.44}_{\pm 0.64}$ $\underline{100.27}_{\pm 0.84}$ $\underline{54.17}_{\pm 0.66}$ $\underline{78.11}_{\pm 0.95}$
HCNN+ ($\mathbb{L}$) $57.69_{\pm 0.52}$ $98.14_{\pm 0.44}$ $52.73_{\pm 0.27}$ $77.98_{\pm 0.32}$
Decrease in % 6.10 2.12 2.66 0.17
@@ 5 Conclusion
In this work, we propose a robust curvature-aware optimization framework for hyperbolic deep learning, addressing challenges in stability, computational efficiency, and overfitting. By introducing Riemannian AdamW, a novel distance rescaling function, and leveraging efficient CUDA implementations, our approach achieved performance improvements in hierarchical metric learning, EEG classification, and image classification tasks, while reducing computational costs, highlighting the practicality of our methods for large-scale applications. These findings emphasize the importance of proper curvature adaptation in hyperbolic learning, paving the way for future research in optimizing hyperbolic embeddings across diverse fields.
6 Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. We are committed to contributing positively to society and human well-being, while respecting the privacy of the subjects involved in our evaluation. Our experiments utilize three established EEG datasets: MI, SSVEP, and ERN, which involve data from human subjects. These datasets are fully anonymized and do not contain any personally identifiable information. Additionally, they are publicly available and widely used within the machine learning community.
Appendix A Operations in hyperbolic geometry
@@
Each Lorentz transformation can be decomposed via polar decomposition into a Lorentz rotation and a Lorentz boost, expressed as $\mathrm{\textbf{A}}=\mathrm{\textbf{R}}\mathrm{\textbf{B}}$ (Moretti, 2002). The rotation matrix R is designed to rotate points around the time axis and is defined as
@@
Appendix B Convolution Trick
To address the computational requirements issue, we adopt an alternative definition of the Lorentz Linear layer from Dai et al. (2021), which decomposes the transformation into a Lorentz boost and a Lorentz rotation. Using this definition, we replace the matrix multiplication employed by Bdeir et al. (2024) for the spatial dimensions and time component projection with a learned rotation operation and a Lorentz boost. Additionally, we can achieve the rotation operation using a parameterization of the convolution weights while still relying on the CUDA convolution implementations, significantly improving computational efficiency.
To apply this concept to the convolution operation, the convolution weights, after unfolding, must form a rotation matrix. We define the dimensions of this matrix as $n=(channels_{in}\cdot kernel_{width}\cdot kernel_{height})$ and $n^{\prime}=channels_{out}$ respectively. We then use either rotation operation presented above to a norm-preserving transformation $\bm{z}=\bm{W^{T}}\bm{x}\cdot\frac{\lVert\bm{x}\rVert}{\lVert\bm{W^{T}}\bm{x}\rVert}$ where $\bm{W}\in\mathbb{R}^{(n^{\prime}\cdot n)}$. This formulation allows us to utilize existing efficient implementations of the convolution operation by directly parameterizing the kernel weights before passing them into the convolutional layer. Finally, we formalize the new Lorentz Convolution as:
$\bm{out}=\text{LorentzBoost(DistanceRescaling(RotationConvolution(}\bm{x}\textbf{)))}$)) (9)
where TanhRescaling is the operation described in Eq.4 and RotationConvolution is a normal convolution parameterized through the procedure in Algorithm 2 where Transform is the norm-preserving transformation above.
Algorithm 2 Lorentz Convolution Parametarization
  $\bm{W}\in\mathbb{R}^{C_{in},C_{out},K_{width},K_{length}}$
  function Adaptweight
     if $K_{width}\cdot K_{length}\cdot C_{in}>C_{out}$ then
        return $\bm{W}$
     else
        $\bm{W}=\bm{W}.reshape(K_{width}\cdot K_{length}\cdot C_{in},C_{out})$
        $\bm{\hat{W}}_{core}=$ Transform ($\bm{W}$)
        return $\bm{\hat{W}}.reshape(C_{in},C_{out},K_{width},K_{length})$
     end if
  end function
Appendix C Scaling Lorentzian Vectors
@@
Appendix D Ablations
D.1 Image Classification
We test the effect of individual model components in Table 6. Each subsequent model involves the default architecture presented in the experimental setup minus the mentioned component. As we can see, the best results are achieved when all the architectural components are included. In the case of attempting to learn the curvature without the proposed optimizer schema, the model breaks completely down due to excessive numerical inaccuracies.
@@ D.1 Image Classification
D.2 Metric Learning
In Table 7 we present the LHIER model results under different optimizers and components.
Table 7: Ablation on components for Metric Learning
Methods Arch. CUB Cars SOP
R@1 R@2 R@1 R@2 R@1 R@10
CNN Backbone
LHIER RSGD R512 64.2 72.1 73.9 81.4 67.8 73.9
LHIER RAdam R512 70.1 79.8 87.6 92.0 79.8 90.3
LHIER CST Curvature R512 71.2 80.8 88.7 91.8 79.1 89.7
LHIER No Learning Schema R512 65.8 72.0 - - - -
LHIER R512 73.4 82.4 90.0 94.0 81.9 93.1
D.3 EEG Classification
We also train the model with the original Adam Optimizer and include the results in Table 8. This clearly shows the improvement provided by our RAdamW optimizer over standard RAdam.
Table 8: Performance comparison for the EEG datasets MI, SSVEP, and ERN. We report the average accuracy for MI and SSVEP and the AUC for ERN. The best result is highlighted in bold.
Models MI SSVEP ERN HyperMAtt + Adam (ours) 68,53$\pm$1.29 62.42$\pm$2.62 74.98$\pm$5.84 HyperMAtt + AdamW (ours) 74.12$\pm$2.91 68.10$\pm$2.41 78.01$\pm$1.30